# Scalable piecewise smoothing with BART

### v3

@@ 5 Experiments
We performed two synthetic data simulations to assess ridgeBART’s ability to perform targeted smoothing (Section 5.1) and to recover more general functions (Section 5.2). We then applied ridgeBART to the NBA shot chart data (Section 5.3).
We ran ridgeBART with $M=50$ trees and $D=1$ ridge functions in each leaf and report the results from using cosine, ReLU, and $\tanh$ activation functions. We compared ridgeBART to the original BART model, Soft BART, GP-BART, and tsBART. We specifically used the implementations available in the BART (Sparapani et al.,, 2021), SoftBart (Linero,, 2022), gpbart (Maia et al., 2024b, ), and tsbart (Starling,, 2024) R packages. For each method, we drew 10,000 posterior samples by sequentially running 10 Markov chains for 2,000 iterations discarding the first 1000 iterations of each chain as burn-in. All experiments were run on a shared high-throughput cluster using R version 4.4.1.
@@ 5.1 Targeted smoothing
Each panel of Figure 2 shows the patient’s observed data (points), true curve $f(\bm{x},z)$ (solid black line), the estimated posterior mean (red line) and pointwise 95% credible intervals (shaded) of $f(\bm{x},z).$ Every method produced excellent estimates of each individual curve: the posterior means closely follow the true curves’ shapes and the pointwise 95% credible intervals display higher-than-nominal coverage. But on further inspection, we see that the BART and Soft BART posterior mean estimates appear somewhat jagged. More concerning, we see that the widths of the credible intervals are quite volatile and oscillate from very narrow to wide. In sharp contrast, the ridgeBART posterior mean estimates are much smoother and there is less variation in the posterior uncertainty across the time domain. Visually, the three ridgeBART estimates appeared much more accurate than the BART and Soft BART estimates.
@@
Figures 3(a) and 3(b) show how each method’s in-sample patient and out-of-sample patient errors compare relative to the original BART model’s as functions of the number of training patients $n.$ For predicting recovery curves for in-sample patients, each ridgeBART model and Soft BART outperformed the original BART model. However, as we increased the number of training individuals $n,$ the gap between BART, Soft BART, and ridgeBART with $\tanh$ activation diminished. In contrast, ridgeBART with cosine and ReLU activations continued to improve relative to the original BART model, achieving 20-40% reductions in RMSE. We additionally observed that ridgeBART with cosine and ReLU activation substantially out-performed BART when forecasting recovery curves for new patients and were competitive with Soft BART.
@@ 5.2 Generalized Smoothing
Although motivated by targeted smoothing, ridgeBART can also be used for generic nonparametric regression. That is, given observations $y\sim\mathcal{N}\left(f(\bm{x}),\sigma^{2}\right),$ we can run ridgeBART with $\bm{z}=\bm{x}$ to approximate $f$ with piecewise-smooth functions. For each $n\in\{500,1000,2000,4000\},$ we generated a dataset $(\bm{x}_{1},y_{1}),\ldots,(\bm{x}_{n},y_{n})$ where each $\bm{x}_{i}$ was drawn uniformly from $[0,1]^{5}$ and $y_{i}\sim\mathcal{N}\left(f(\bm{x}_{i}),1\right)$ where $f(\bm{x})=\sin(\pi x_{1}x_{2})+20(x_{3}-0.5)^{2}+10x_{4}+5x_{5}$ is the Friedman function frequently used to benchmark BART extensions.
@@ 5.3 Shot Charts
Returning to our motivating example, recall that our goal is to estimate the field goal percentage (FG%) of every NBA player at every location on the court. We obtained a dataset of 196,095 shots from the 2023-2024 season, including the playoffs, using the hoopR package (Gilani,, 2023) consisting of a binary indicator $y$ of whether a shot was made ($y=1$) or missed ($y=0$); the horizontal and vertical coordinates of the shot ($\bm{z}$); and a vector of covariates $\bm{x}$ about the player who attempted the shot. In addition to the categorical player identity, $\bm{x}$ included their position (guard, forward, or center), height, and weight. By including these covariates into our model, we can attempt to “borrow strength” across players at the same position and with similar physical characteristics.
Using 20-fold cross-validation, we compared the predictive quality of ridgeBART models fit with different activation functions to three alternative models. First, we fit two generalized additive models (GAM) to these data using the mgcv package (Wood,, 2017). The first pooled together data from all players, estimating FGas a function of location alone. The second model fits separate GAM models to the data for each player. Interestingly, we found that the completely pooled GAM performed substantially better out-of-sample than the completely un-pooled, player-specific model. We additionally compare the ridgeBART and GAM fits to that produced by BART.
@@
Figure 4 compares the FGheat maps estimated by ridgeBART with cosine activation for Stephen Curry, Ben Simmons, Giannis Antetokounmpo, and LeBron James. These players are known to score in different ways. Simmons and Antetokounmpo are known for being physical and scoring at close-range, while Curry is instead known for being a long-range shooter. James, in contrast, is considered to be a good all-around scorer. The comparisons between these players are illuminated by the ridgeBART shot charts, all of which do not have the same pathological sharp discontinuities as the BART shot charts. Clearly, we can see that Stephen Curry is a significantly better shooter at all areas of the court compared to James, Antetokounmpo, and Simmons. James and Antetokounmpo have very similar looking shot charts, where the main difference is that Antetokounmpo is slightly better in the immediate vicinity of the basket. These results suggest that, despite their differences in play-style, James and Antetokounmpo are equally-skilled at shooting the basketball.
@@ C.1 Ridge approximations anisotropic Hölder class functions
In each subdomain $\Psi_{r}\subseteq[0,1]^{p}$, if the function $f_{0}\in\mathcal{H}_{\lambda}^{\mathcal{A},p}$, then there exists a ridge function approximation
$f_{D}^{r}(\bm{x})=\sum_{l=1}^{D}\beta_{l}^{(r)}\varphi(\omega_{l}^{(r)\top}\bm{x}+b_{l}^{(r)}),$
such that
@@
For each dimension $j$, we define the partition
$10=x_{j,0}<x_{j,1}<\dots<x_{j,N_{j}}=$ 1
where $N_{j}$ is an integer chosen such that $N_{j}\approx D^{1/p}$. Thus, each sub-interval has length $\Delta_{j}=1/N_{j}$. We denote the sub-boxes by
$B_{\bm{k}}=[x_{1,k_{1}},x_{1,k_{1}+1}]\times\dots\times[x_{p,k_{p}},x_{p,k_{p}+1}],\ \ \bm{k}=(k_{1},\dots,k_{p}).$
Following the ideas of piecewise polynomial approximation in Barron, (1993) and Pinkus, (1997), in each box $B_{\bm{k}}$, we define $P_{\bm{k}}(\bm{x})$ to be the unique multilinear polynomial that matches $f_{0}$ at the $2^{p}$ corners. That is, if $\bm{v}$ is a corner, then $P_{\bm{k}}(\bm{v})=f_{0}(\bm{v})$. This leads us to define the global piecewise polynomial
$P^{r}(\bm{x})\coloneqq\sum_{\bm{k}}[P_{\bm{k}}(\bm{x})\mathbbm{1}(\bm{x}\in B_{\bm{k}})].$
Again, note that for any $\bm{x}\in B_{\bm{k}}$, if we pick a corner $\bm{v}$, since $P_{\bm{k}}(\bm{v})=f_{0}(\bm{v})$, we have
$|f_{0}(\bm{x})-P_{\bm{k}}(\bm{x})|\leq|f_{0}(\bm{x})-f_{0}(\bm{v})|$
But $\|\bm{x}-\bm{v}\|\leq\max_{j}\Delta_{j}$. Now, in order to upper bound $|f_{0}(\bm{x})-f_{0}(\bm{v})|$, we use the repeated integration argument for a single coordinate following the lines of Elliott, (1994). Fix a coordinate $j$. We define a path that changes $\bm{v}$ to $\bm{x}$ only in the $j$-th coordinate, leaving other coordinates the same. Concretely, define
$\bm{x}_{j}(t)\coloneqq(v_{1},\dots,v_{j-1},t,v_{j+1},\dots,v_{p}),\ t\in[\min(v_{j},x_{j}),\max(v_{j},x_{j})]$
So $\bm{x}_{j}(\cdot)$ is a straight line in the $j$-th coordinate from $v_{j}$ to $x_{j}$, keeping the others fixed at $v_{i}$. Now, define the univariate function $\tilde{f}_{j}(t)\coloneqq f_{0}(\bm{x}_{j}(t))$. Since $f_{0}$ is $\lfloor\alpha_{r,j}\rfloor$ times differentiable in $x_{j}$ (holding others fixed), $\tilde{f}_{j}(t)$ is $\lfloor\alpha_{r,j}\rfloor$ times differentiable in $t$. Thus,
$f_{0}(\bm{x})-f_{0}(\bm{v})=f_{0}(\bm{x}_{j}(x_{j}))-f_{0}(\bm{x}_{j}(v_{j}))=\tilde{f}_{j}(x_{j})-\tilde{f}_{j}(v_{j})$
The repeated integral argument is elaborated below. Let $m_{j}=\lfloor\alpha_{r,j}\rfloor$ and $\tau_{j}=\{\alpha_{r,j}\}\equiv\alpha_{r,j}-m_{j}$. Then:
[…]
Again, if $m_{j}\geq 2$, we can write,
[…]
Then
[…]
Proceeding to the $k^{th}$ derivative, eventually we reach:
[…]
By integrating $m_{j}$ times, we accumulate polynomial factors in $(u-v_{j})$ of integer degrees, plus a “final leftover” that depends on $\tilde{f}_{j}^{(m_{j})}(\cdot)$. From Definition 1, we have:
$|\tilde{f}_{j}^{(m_{j})}(u)-\tilde{f}_{j}^{(m_{j})}(v_{j})|\leq\lambda|u-v_{j}|^{\tau_{j}}.$
Integrating from $v_{j}$ to $x_{j}$ yields a remainder of order $(x_{j}-v_{j})^{m_{j}+\tau_{j}}=(x_{j}-v_{j})^{\alpha_{r,j}}$. The standard bounding is
[…]
Summing all contributions,
$|\tilde{f}_{j}(x_{j})-\tilde{f}_{j}(v_{j})|\lesssim|x_{j}-v_{j}|^{\alpha_{r,j}}.$
We now do this for each coordinate $j=1,\dots,p$, changing from $v_{j}$ to $x_{j}$. Concretely, define a path in $\mathbb{R}^{p}$ that changes $\bm{v}$ into $\bm{x}$ one coordinate at a time. Denote $\bm{z}_{0}=\bm{v}$, $\bm{z}^{(l+1)}=\bm{z}^{(l)}+\Delta_{j(l)}$, where $j(\ell)$ indicates which coordinate is being changed at step $\ell$. Summing the difference in each step, each dimension $j$ contributes a factor $(x_{j}-v_{j})^{\alpha_{r,j}}$. We thus obtain:
$|f_{0}(\bm{x})-f_{0}(\bm{v})|\lesssim\sum_{j=1}^{p}(\Delta_{j})^{\alpha_{r,j}}.$
This implies that
$\sup_{\bm{x}\in B_{\bm{k}}}|f_{0}(\bm{x})-P_{\bm{k}}(\bm{x})|\lesssim\sum_{j=1}^{p}N_{j}^{-\alpha_{r,j}}$
Combining the sub-boxes, we have:
$\sup_{\bm{x}\in\Psi_{r}}|f_{0}(\bm{x})-P^{r}(\bm{x})|$ $=\max_{\bm{k}}\sup_{\bm{x}\in B_{\bm{k}}}|f_{0}(\bm{x})-P_{\bm{k}}(\bm{x})|$
$\lesssim\sum_{j}N_{j}^{-\alpha_{r,j}}$
$\leq CD^{-\min_{j}\alpha_{r,j}/p}$ (C1)
where $C>0$ is an appropriate constant. Now, following the lines of tensor-product spline construction in Yarotsky, (2017), we can encode each polynomial patch $P_{\bm{k}}$ as a finite linear combination of univariate basis functions, that is,
$P_{\bm{k}}(\bm{x})=\sum_{k_{1}=0}^{m}\cdots\sum_{k_{p}=0}^{m}c_{k_{1},\dots k_{p}}\psi_{k_{1}}(x_{1})\cdots\psi_{k_{p}}(x_{p})$
A standard argument in Yarotsky, (2017) and Barron, (1993) states that each univariate $\psi_{k_{j}}(x_{j})$ can be exactly or approximately encoded as a sum of ridge units $\varphi(\bm{e}_{j}^{\top}\bm{x}+\bm{b})$, (i.e. a “ridge” that depends on only the $j^{th}$ coordinate). Then their product is expanded into a sum of such units. Summing over all the multi-indices $(\bm{k}_{1},\cdots,\bm{k}_{p})$ yields a single sum-of-ridge representation. Thus, each local polynomial $P_{\bm{k}}$ is replace by a sum-of-ridge function $\tilde{P}_{\bm{k}}$ with at most a constant factor blow-up in the cardinality. Summing across all sub-boxes, we define
$f_{D}^{r}(\bm{x})=\sum_{k}[\tilde{P}_{\bm{k}}(\bm{x})\mathbbm{1}(\bm{x}\in B_{\bm{k}})]\implies\|P^{r}-f^{r}_{D}\|_{\infty}\leq\epsilon$ (C2)
for any small $\epsilon>0$, if $\varphi$ is suitably universal. One can take $\epsilon\approx 0$ if for instance, $\varphi$ is piecewise polynomial (like ReLU). Now, collating the bounds from Sections C.1 and C2, we have by the triangle inequality:
$\|f_{0}(\bm{x})-f^{r}_{D}(\bm{x})\|_{\infty}$ $\leq\|f_{0}(\bm{x})-P^{r}(\bm{x})\|_{\infty}+\|P^{r}(\bm{x})-f^{r}_{D}(\bm{x})\|_{\infty}$
$\leq CD^{-\min_{j}\alpha_{r,j}/p}+\epsilon.$
Choosing $\epsilon\approx D^{-\min_{j}\alpha_{r,j}/p}$ yields the desired local ridge approximation result. ∎
To recall, the goal is to approximate any function $f_{0}\in\mathcal{H}_{\lambda}^{\mathcal{A},p}$ globally using a ridge function approximation of the form
$\hat{f}(\bm{x})=\sum_{d=1}^{D^{\star}}\beta_{d}\varphi(\bm{\omega}_{d}^{\top}\bm{x}+b_{d})$
where we ensure that $D^{\star}\approx RD$.
Now, we proceed to prove Lemma 1.

### v4

@@ 5 Experiments
We performed two synthetic data simulations to assess ridgeBART’s ability to perform targeted smoothing (Section 5.1) and to recover more general functions (Section 5.2). We then applied ridgeBART to the NBA shot chart data (Section 5.3). We ran ridgeBART with $M=50$ trees and $D=1$ ridge functions in each leaf and report the results from using cosine, ReLU, and $\tanh$ activation functions. We compared ridgeBART to the original BART model, Soft BART, GP-BART, and tsBART. We specifically used the implementations available in the BART (Sparapani et al.,, 2021), SoftBart (Linero,, 2022), gpbart (Maia et al., 2024b, ), and tsbart (Starling,, 2024) R packages. For each method, we drew 10,000 posterior samples by sequentially running 10 Markov chains for 2,000 iterations discarding the first 1000 iterations of each chain as burn-in. All experiments were run on a shared high-throughput cluster using R version 4.4.1.
@@ 5.1 Targeted smoothing
Each panel of Figure 2 shows the patient’s observed data (points), true curve $f(\bm{x},z)$ (solid black line), the estimated posterior mean (red line) and pointwise 95% credible intervals (shaded) of $f(\bm{x},z).$ Every method produced excellent estimates of each individual curve: the posterior means closely follow the true curves’ shapes and the pointwise 95% credible intervals display higher-than-nominal coverage. But on further inspection, we found that BART and SoftBART returned somewhat jagged posterior mean estimates and credible intervals whose widths oscillated from very narrow to very wide. In sharp contrast, ridgeBART returned much smoother posterior means and uncertainty intervals with more consistent width. Visually, the three ridgeBART estimates also appeared much more accurate than the BART and SoftBART estimates.
@@
Figures 3(a) and 3(b) show how each method’s in-sample patient and out-of-sample patient errors compare to the original BART model’s as functions of the number of training patients $n.$ For predicting recovery curves for in-sample patients, each ridgeBART model and SoftBART outperformed the original BART model. However, as we increased the number of training individuals $n,$ the gap between BART, SoftBART, and ridgeBART with $\tanh$ activation diminished. In contrast, ridgeBART with cosine and ReLU activations continued to improve relative to the original BART model, achieving 20-40% reductions in RMSE. We additionally observed that ridgeBART with cosine and ReLU activation substantially out-performed BART when forecasting recovery curves for new patients and were competitive with SoftBART.
@@ 5.2 Generalized Smoothing
Although motivated by targeted smoothing, ridgeBART can also be used for generic nonparametric regression. That is, given observations $y\sim\mathcal{N}\left(f(\bm{x}),\sigma^{2}\right),$ we can run ridgeBART with $\bm{z}=\bm{x}$ to approximate $f$ with piecewise-smooth functions. For each $n\in\{500,1000,2000,4000\},$ we generated a dataset $(\bm{x}_{1},y_{1}),\ldots,(\bm{x}_{n},y_{n})$ where each $\bm{x}_{i}$ was drawn uniformly from $0,1]^{5}$ and $y_{i}\sim\mathcal{N}\left(f(\bm{x}_{i}),1\right)$ where $f(\bm{x})=\sin(\pi x_{1}x_{2})+20(x_{3}-0.5)^{2}+10x_{4}+5x_{5}$ is the Friedman function frequently used to benchmark BART extensions.
@@ 5.3 Shot Charts
Returning to our motivating example, recall that our goal is to estimate the field goal percentage (FG%) of every NBA player at every location on the court. We obtained a dataset of 196,095 shots from the 2023-2024 season, including the playoffs, using the hoopR package (Gilani,, 2023) consisting of a binary indicator $y$ of whether a shot was made ($y=1$) or missed ($y=0$); the horizontal and vertical coordinates of the shot ($\bm{z}$); and a vector of covariates $\bm{x}$ about the player who attempted the shot. In addition to the categorical player identity, $\bm{x}$ included their position (guard, forward, or center), height, and weight. By including these covariates into our model, we can “borrow strength” position and physical characteristics.
Using 20-fold cross-validation, we compared the predictive quality of ridgeBART models fit with different activation functions to three alternative models. First, we fit two generalized additive models (GAM) to these data using the mgcv package (Wood,, 2017). The first pooled together data from all players, estimating FG% as a function of location alone. The second model fits separate player-specific GAMs. Interestingly, the completely pooled model performed substantially better out-of-sample than the un-pooled, player-specific models. We additionally compare the ridgeBART and GAM fits to that produced by BART.
@@
Figure 4 compares the FG% heat maps estimated by ridgeBART with cosine activation for Stephen Curry, Ben Simmons, Giannis Antetokounmpo, and LeBron James. These players are known to score in different ways. Simmons and Antetokounmpo are known for being physical and scoring at close-range, while Curry is instead known for being a long-range shooter. James, in contrast, is considered to be a good all-around scorer. The comparisons between these players are illuminated by the ridgeBART shot charts, all of which do not have the same pathological sharp discontinuities as the BART shot charts. Clearly, we can see that Stephen Curry is a significantly better shooter at all areas of the court compared to James, Antetokounmpo, and Simmons. James and Antetokounmpo have very similar looking shot charts, where the main difference is that Antetokounmpo is slightly better in the immediate vicinity of the basket. These results suggest that, despite their differences in play-style, James and Antetokounmpo are equally-skilled at shooting the basketball.
@@ C.1 Ridge approximations anisotropic Hölder class functions
Let $f_{0}\in\mathcal{H}_{\lambda}^{\mathcal{A},p}$ be a function in the piecewise anisotropic Hölder class. Then for each subdomain $\Psi_{r}\subseteq[0,1]^{p}$, there exists a ridge function approximation $f_{D}^{r}(\bm{x})=\sum_{l=1}^{D}\beta_{l}^{(r)}\varphi(\bm{\omega}_{l}^{(r)\top}\bm{x}+b_{l}^{(r)})$ such that
@@
for some constant $C_{r}$ that depends on $\lambda$ and $\bm{\alpha}_{r}$.
@@
We first establish a rigorous bound on the modulus of continuity for $f_{0}$ based on its Hölder properties.
(1) Bounding the modulus of continuity. Let $\bm{x},\bm{v}$ be two points in the subdomain $\Psi_{r}$. We wish to bound $|f_{0}(\bm{x})-f_{0}(\bm{v})|$. We can construct a path from $\bm{v}$ to $\bm{x}$ that changes one coordinate at a time. Let $\bm{z}_{0}=\bm{v}$, $\bm{z}_{j}=(x_{1},\dots,x_{j},v_{j+1},\dots,v_{p})$, and $\bm{z}_{p}=\bm{x}$. Then, by the triangle inequality:
$|f_{0}(\bm{x})-f_{0}(\bm{v})|=|f_{0}(\bm{z}_{p})-f_{0}(\bm{z}_{0})|\leq\sum_{j=1}^{p}|f_{0}(\bm{z}_{j})-f_{0}(\bm{z}_{j-1})|$
Each term $|f_{0}(\bm{z}_{j})-f_{0}(\bm{z}_{j-1})|$ involves a change only in the $j$-th coordinate. Let $m_{j}=\lfloor\alpha_{r,j}\rfloor$ and $\tau_{j}=\alpha_{r,j}-m_{j}$. For the univariate slice function $\tilde{f}_{j}(t)=f_{0}(x_{1},\dots,x_{j-1},t,v_{j+1},\dots,v_{p})$, Taylor’s theorem with integral remainder gives:
$\tilde{f}_{j}(x_{j})-\tilde{f}_{j}(v_{j})=\sum_{k=1}^{m_{j}-1}\frac{\tilde{f}_{j}^{(k)}(v_{j})}{k!}(x_{j}-v_{j})^{k}+\int_{v_{j}}^{x_{j}}\frac{(x_{j}-t)^{m_{j}-1}}{(m_{j}-1)!}\tilde{f}_{j}^{(m_{j})}(t)dt$
The Hölder condition from Definition 1 states $|\tilde{f}_{j}^{(m_{j})}(t)-\tilde{f}_{j}^{(m_{j})}(v_{j})|\leq\lambda|t-v_{j}|^{\tau_{j}}$. Using this, the difference $|f_{0}(\bm{z}_{j})-f_{0}(\bm{z}_{j-1})|=|\tilde{f}_{j}(x_{j})-\tilde{f}_{j}(v_{j})|$ can be shown to be bounded by a constant times $|x_{j}-v_{j}|^{\alpha_{r,j}}$. Summing these contributions gives the bound on the modulus of continuity:
$|f_{0}(\bm{x})-f_{0}(\bm{v})|\leq C_{1}\sum_{j=1}^{p}|x_{j}-v_{j}|^{\alpha_{r,j}}$
for a constant $C_{1}$ depending on $\lambda$ and $\bm{\alpha}_{r}$. This is the key property we will use. We then use this result to bound the error of approximating $f_{0}$ with a specially constructed multivariate polynomial, $P_{N}(\bm{x})$.
(2) Polynomial approximation via a kernel operator. We construct a polynomial approximant for $f_{0}$ using a convolution with a polynomial kernel. Let $K_{N}(t)$ be the univariate Jackson kernel, a non-negative polynomial of degree less than $2N$ with specific properties (see Jordõ and Manegatto, (2014) for an extensive review). We define a multivariate tensor-product kernel $\mathcal{K}_{N}(\bm{z})=\prod_{j=1}^{p}K_{N}(z_{j})$. We can now define a polynomial approximation operator $\mathcal{P}_{N}$ as:
$P_{N}(\bm{x})=(\mathcal{P}_{N}f_{0})(\bm{x})=\int_{\mathbb{R}^{p}}f_{0}(\bm{y})\mathcal{K}_{N}(\bm{x}-\bm{y})d\bm{y}$
By a change of variables $\bm{z}=\bm{x}-\bm{y}$, this becomes:
$P_{N}(\bm{x})=\int_{\mathbb{R}^{p}}f_{0}(\bm{x}-\bm{z})\mathcal{K}_{N}(\bm{z})d\bm{z}$
The Jackson kernel is constructed such that $\int K_{N}(t)dt=1$, which implies $\int\mathcal{K}_{N}(\bm{z})d\bm{z}=1$. We can now bound the approximation error:
$|f_{0}(\bm{x})-P_{N}(\bm{x})|$ $=\left|f_{0}(\bm{x})\cdot 1-\int f_{0}(\bm{x}-\bm{z})\mathcal{K}_{N}(\bm{z})d\bm{z}\right|$
$=\left|\int[f_{0}(\bm{x})-f_{0}(\bm{x}-\bm{z})]\mathcal{K}_{N}(\bm{z})d\bm{z}\right|$
$\leq\int|f_{0}(\bm{x})-f_{0}(\bm{x}-\bm{z})|\mathcal{K}_{N}(\bm{z})d\bm{z}$
Now, we use the modulus of continuity bound from Step 1:
$|f_{0}(\bm{x})-P_{N}(\bm{x})|\leq\int\left(C_{1}\sum_{j=1}^{p}|z_{j}|^{\alpha_{r,j}}\right)\mathcal{K}_{N}(\bm{z})d\bm{z}=C_{1}\sum_{j=1}^{p}\int|z_{j}|^{\alpha_{r,j}}\mathcal{K}_{N}(\bm{z})d\bm{z}$
Since the kernel $\mathcal{K}_{N}(\bm{z})$ is a product of univariate kernels, the integral separates:
$\int|z_{j}|^{\alpha_{r,j}}\mathcal{K}_{N}(\bm{z})d\bm{z}=\left(\int|z_{j}|^{\alpha_{r,j}}K_{N}(z_{j})dz_{j}\right)\prod_{i\neq j}\left(\int K_{N}(z_{i})dz_{i}\right)=\int_{-\infty}^{\infty}|t|^{\alpha_{r,j}}K_{N}(t)dt$
A key property of the Jackson kernel is that for any $\alpha>0$, $\int|t|^{\alpha}K_{N}(t)dt\leq C_{\alpha}N^{-\alpha}$. Applying this, we get:
$|f_{0}(\bm{x})-P_{N}(\bm{x})|\leq C_{1}\sum_{j=1}^{p}C_{\alpha_{r,j}}N^{-\alpha_{r,j}}\leq C_{2}N^{-\min_{j}\alpha_{r,j}}$
The polynomial $P_{N}(\bm{x})$ has degree less than $2N$ in each coordinate and thus has $O(N^{p})$ coefficients. We link this to our ridge function budget $D$ by setting $D\asymp N^{p}$, which implies $N\asymp D^{1/p}$. Substituting this into the error bound gives:
$\sup_{\bm{x}\in\Psi_{r}}|f_{0}(\bm{x})-P_{N}(\bm{x})|\leq C_{2}(D^{1/p})^{-\min_{j}\alpha_{r,j}}=C_{2}D^{-\min_{j}\alpha_{r,j}/p}$ (C1)
(3) Ridge function approximation of the polynomial. Now we must show that the polynomial $P_{N}(\bm{x})$ from the previous step can be well-approximated by our target structure, a sum of $D$ ridge functions, $f_{D}^{r}(\bm{x})$. While foundational universal approximation theorems (Leshno et al.,, 1993; Cybenko,, 1989) guarantee the existence of such an approximation, they do not provide the explicit rate of convergence needed for our argument.
To establish a rate, we leverage quantitative results from approximation theory for neural networks, which show that networks can approximate smooth functions with extreme efficiency. The key insight is that our intermediate function, the polynomial $P_{N}(\bm{x})$, is infinitely differentiable ($C^{\infty}$) on the compact subdomain $\Psi_{r}$. This high degree of smoothness allows for an approximation rate that is much faster than the rate for functions that are merely Hölder continuous.
According to established results in approximation theory (see, e.g., Pinkus, (1997, Theorem 6.8)), a function residing in a Sobolev space $W^{k,\infty}(\Psi_{r})$ can be approximated by a shallow network (a sum of ridge functions) with $D$ neurons with an error bounded by $O(D^{-k/p})$. Since our polynomial $P_{N}$ is in $W^{k,\infty}(\Psi_{r})$ for any arbitrarily large integer $k$, we can select $k$ such that the approximation error is negligible compared to the primary error from the previous step.
Let’s choose an integer $k$ large enough so that $k>\min_{j}\alpha_{r,j}$. For this $k$, the theory guarantees there exists a ridge function approximant $f_{D}^{r}(\bm{x})=\sum_{l=1}^{D}{\beta_{l}^{(r)}\varphi(\bm{\omega}_{l}^{(r)\top}\bm{x}+b_{l}^{(r)})}$ with $D$ neurons such that:
$\sup_{\bm{x}\in\Psi_{r}}|P_{N}(\bm{x})-f_{D}^{r}(\bm{x})|\leq C_{3}D^{-k/p}$ (C2)
Because we chose $k>\min_{j}\alpha_{r,j}$, the rate $D^{-k/p}$ is asymptotically much faster than the rate $D^{-\min_{j}\alpha_{r,j}/p}$ from the polynomial approximation step.
Finally, we use the triangle inequality to combine the two error bounds from (C1) and (C2):
$\sup_{\bm{x}\in\Psi_{r}}|f_{0}(\bm{x})-f_{D}^{r}(\bm{x})|$ $\leq\sup_{\bm{x}\in\Psi_{r}}|f_{0}(\bm{x})-P_{N}(\bm{x})|+\sup_{\bm{x}\in\Psi_{r}}|P_{N}(\bm{x})-f_{D}^{r}(\bm{x})|$
$\leq C_{2}D^{-\min_{j}\alpha_{r,j}/p}+C_{3}D^{-\min_{j}\alpha_{r,j}/p}$
$\leq C_{r}D^{-\min_{j}\alpha_{r,j}/p}$
where $C_{r}=C_{2}+C_{3}$ is a constant depending on $\lambda$ and $\bm{\alpha}_{r}$. This completes the proof. ∎